OpenStudy (anonymous):

Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart? I know its 3 hours, but i dont know how to find it the proper way.

OpenStudy (anonymous):

D=RT 45=(10+5)T 45=15T T=445/15 T=3 HOURS ANS.